### v1

# Token-Importance Guided Direct Preference Optimization

Abstract
Ensuring that large language models (LLMs) generate outputs aligned with human preferences is important for safe and effective AI interactions. While Direct Preference Optimization (DPO) employs an implicit reward function to optimize the policy model, however, it and its related variants overlook the differential importance of individual tokens and are sensitive to judgment noise in preference datasets during generation. Although recent methods attempt to assess the important weight of tokens via probability prediction or simplistic weighting schemes, these evaluation methods are prone to biases and still cannot fully address these issues. To solve this problem, we propose the Token-Importance Guided Direct Preference Optimization (TI-DPO), which introduces two key innovations: the gradient-based token-importance weights that dynamically prioritize critical tokens, and a triple loss that explicitly guides model outputs to approach human-preferred responses and stay away from non-preferred responses. Experimental results show that TI-DPO achieves higher accuracy and stronger generative diversity, providing more stable and computationally efficient solutions compared with DPO and other RLHF methods.
1 Introduction
LLM has shown proficiency in Natural Language Processing (NLP) [1, 2], logical reasoning [3], and code generation [4], emerging as a focal point of recent research. However, as models may generate outputs inconsistent with intended purposes or ethical standards, human preference alignment aims to ensure that LLMs adhere to human values [5, 6], producing beneficial and harmless content. Against this backdrop, Reinforcement Learning from Human Feedback (RLHF) has become a prevailing approach for achieving alignment [7, 8, 9]. It leverages human-annotated preference data to train reward models and fine-tunes LLMs using Reinforcement Learning (RL) methods [10, 11] like Proximal Policy Optimization (PPO) [12]. While effective, RLHF suffers from inherent limitations: its reliance on iterative policy optimization and complex reward modeling leads to training instability and exorbitant computational costs [13].
Addressing these constraints, Direct Preference Optimization (DPO) [14] eliminates the necessity for a reward model. Instead, it employs direct pairwise preference comparisons to optimise model policies. Inspired by DPO’s implicit reward mechanism, a series of preference optimization models have been proposed in recent years, such as ORPO [7], f-DPO [15] and CPO [16]. However, DPO’s loss function does not account for the varying importance of individual tokens during the generation of preferred responses [17], further intensifying the problem of sampling distribution change. Consider a model generating medical advice: a single misweighted token, for example, "safe" versus "risk", could drastically alter the response’s safety. Traditional DPO treats all tokens equally, potentially overlooking such critical distinctions.
Motivated by these challenges, researchers have proposed token-level variants of DPO such as TDPO [17], Logic-RL [3], RTO [18] and SimPO [19], aiming to decompose preference alignment into fine-grained token contributions. Based on these methods, some works have additionally considered the impact of the token importance weights, but they often rely on probabilistic proxies [20] or simplistic weighting schemes [21], which are prone to bias and fail to fully leverage the structural relationships between semantically similar responses. By contrast, we propose Token-Importance Guided Direct Preference Optimization (TI-DPO), using gradient attribution to dynamically compute token-importance weights. Token importance weights act as a ‘spotlight’, focusing optimization on tokens that most influence human judgments.
Furthermore, recent studies [20, 19] have noticed some shortcomings in DPO’s design: it compares merely "good" versus "bad" samples. Concurrently, LLMs generate preference data through various methods [22], such as rejection sampling [23] and employing different prompts for distinct preference levels [22], which spurred us to define different levels of samples. Under such a background, our TI-DPO method constructs a "good-bad" triplet relationship by incorporating the intermediate generated outputs [24]. This triplet structure explicitly guides the intermediate output to approach human preferences and distance from non-preferred responses, achieving fine-grained preference alignment and promoting a continuous gradient of preference learning, rather than relying solely on binary comparisons.
The following contributions are made in the course of this work:
  • Based on DPO, we propose TI-DPO, a token-level approach with important weights. Integrating triplet loss with gradient-based token importance weighting, TI-DPO focuses optimization on critical tokens, mitigating distribution shift and improving alignment precision.
  • Theoretically, we prove TI-DPO achieves a tighter loss bound than DPO (Theorem 4.2), ensuring more stable optimization. This theorem formally provides a new perspective on comprehending the superiority of TI-DPO in terms of alignment accuracy.
  • Experiment results indicate that TI-DPO surpasses existing methods in aligning LLMs with human preferences. In preference comparisons, TI-DPO achieves the highest accuracy on TruthfulQA and IFEval tasks, and we compare TI-DPO with three base instruction models across multiple task dimensions with radar graphs. Additionally, ablation experiments confirm the necessity of its token-importance guidance and triplet loss components.
2 Related Work
Human Preference Alignment Human preference alignment has emerged as a critical research paradigm in recent years, focusing on enabling model responses to align with human values and preferences. Early advancements mainly focused on RLHF [26, 27] based on PPO [12]. Due to PPO consuming substantial training resources, some recently proposed RL methods aim to reduce overhead by decreasing parameters of PPO [13] or abandoning the critic model [28]. However, these RL methods may suffer from overfitting in optimal responses. To mitigate this issue, Hu et al. [8] introduced the Reinforce++ model, which employs batch-wise standardized rewards to prevent overfitting and enhance the prompt diversity during training. Concurrently, beyond RL approaches, Rafailov et al. introduced DPO [14], which obviates the need for explicit reward modeling through implicit preference learning. This implicit reward mechanism has inspired a wave of subsequent works [29], such as the SimPO algorithm proposed by Meng et al. [19], which utilizes the sequence-averaged log probability as an implicit reward signal to streamline optimization. Notwithstanding these advancements, DPO’s reliance on large-scale human-annotated preference datasets [30] has motivated derivative studies [31, 9] aimed at reducing data requirements. A notable example is RS-DPO [32], which integrates rejection sampling (RS) with DPO to alleviate data scarcity. However, a more fundamental limitation pertains to the binary nature of traditional preference labels. Although the KTO method proposed by Ethayarajh et al. [33] effectively reduces the reliance on paired preference labels in DPO, most current RLHF and related preference optimizations still mainly rely on binary comparisons between "good" and "bad" responses [34, 7]. Such coarse-grained supervision has obvious shortcomings: human preferences often show continuous gradient differences rather than simply "good" and "bad". Against this backdrop, the triplet optimization method we propose can achieve fine-grained preference alignment.
From Sequence-Level to Token-Level Since RLHF and DPO are effective sequence-level training techniques, KL divergence and reward models are computed at the sequence [16] or sentence level [35], overlooking differences in token importance, which leads to a significant decline in diversity [17]. These limitations have spurred researchers’ research into step [36] or token-level [37, 29] alignment mechanisms, seeking to address the granularity mismatch between coarse-grained sequence rewards and fine-grained token contributions [38]. Lai et al. [39] initiated this research path through Step-DPO, which breaks down the reasoning process into a series of sequential decision points, rather than accessing the entire answer comprehensively. In [17], to address the significant decline in model generation diversity, Token-level DPO (TDPO) reanalyzed and optimized the entire alignment process from the token-level perspective. An additional limitation of RLHF and DPO lies in the fact that rewards are only assigned to the final token, with all other tokens receiving no learning rewards [18]. Meanwhile, Xie et al. [3] proposed a correlation between the frequency of certain tokens and model performance, which inspires us to consider reassigning token weights. In a related vein, Liu et al. proposed Token-level Importance Sampling for DPO (TIS-DPO) [20], which estimates token importance weights using prediction probability differences. Nevertheless, this probabilistic weighting scheme is prone to bias when contrastive models produce inconsistent outputs or fail to capture subtle semantic nuances of human preferences. In contrast, our approach employs gradient-based importance estimation [40], directly quantifying each token’s causal contribution to reward prediction. By focusing on actual gradient impacts rather than probabilistic proxies, this method enhances the precision of aligning model outputs with human preferences.
3 Preliminaries
Before formally elaborating the TI-DPO method, this section first introduces relevant preparatory knowledge to lay the foundation for the subsequent theoretical derivation and model construction.
3.1 Human Preference Alignment
[…]
${\mathcal{L}_{\text{base}}=-\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}\left(x,y_{\text{w}}\right)-r_{\phi}\left(x,y_{\text{l}}\right)\right)\right]}.$ (1)
Here $\sigma(\cdot)$ is sigmoid activation function. Reward model evaluates the LLM’s responses by applying the BT model for ranking losses [26]:
${p\left(y_{\text{w}}\succ y_{\text{l}}\mid x\right)=\frac{\exp\left(r_{\phi}\left(x,y_{\text{w}}\right)\right)}{\exp\left(r_{\phi}\left(x,y_{\text{w}}\right)\right)+\exp\left(r_{\phi}\left(x,y_{\text{l}}\right)\right)}},$ (2)
Nevertheless, the RLHF approach has been acknowledged to encounter challenges related to implementation complexity and training instability as highlighted in a study [42]. DPO [14] derived the optimal policy $\pi^{*}_{\theta}$ from the reward model:
$\pi^{*}_{\theta}(y|x)=\frac{1}{Z(x)}\pi_{\text{ref}}(y\mid x)\text{exp}(\frac{1}{\beta}r_{\phi}(x,y)),$ (3)
In Eq.3, $\pi^{*}_{\theta}$ represents the optimal policy, with $\pi_{\text{ref}}$ denoting the initial model. The partition function $Z(x)$ serves to normalize the policy’s probability distribution[14]. The parameter $\beta$ regulates the extent of divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$. DPO rearranges this equation to express the reward as $r_{\phi}(x,y)=\beta\log\frac{\pi_{\theta}^{*}(y|x)}{\pi_{\text{ref}}(y|x)}-\log Z(x)$. Let the input prompt be represented as $x=[x_{1},x_{2},\ldots,x_{m}]$ and the first $t-1$ tokens generated by the model be denoted as $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. Let $T_{\text{w}}$ and $T_{\text{l}}$ denote the number of preferred tokens and less preferred tokens, respectively. The token-level DPO optimization objective is given by
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}\left(y_{\text{w}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{w}}|x\right)}-\log\frac{\pi_{\theta}\left(y_{\text{l}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{l}}|x\right)}\right)\right)\right],$ (4)
3.2 Triplet Loss
Triplet loss, a powerful loss function for learning embeddings, ensures that within the embedding space, an anchor input is closer to positive inputs than to negative ones. This mechanism enhances the model’s capacity to differentiate between data points that are more or less similar. By simultaneously learning from the similarities and differences among sampled data points, the model is better aligned with human evaluations. The triplet loss operates with triplets $(x_{i},x_{j},x_{k})$, and is designed such that the representation of the anchor $x_{i}$ is nearer to a similar data point $x_{j}$ than to a dissimilar one $x_{k}$. This targeted learning strategy is instrumental in sharpening the model’s feature discrimination, thereby improving its ability to make decisions that resonate with human preferences. The triplet loss is given by
$\mathcal{L}_{\operatorname{trp}}=\sum_{i,j,k}^{T}\left[\left\|f\left(x_{i}\right)-f\left(x_{j}\right)\right\|^{2}_{2}-\left\|f\left(x_{i}\right)-f\left(x_{k}\right)\right\|^{2}_{2}+\alpha_{\operatorname{trp}}\right]_{+}.$ (5)
Here $[z]_{+}$ denotes the rectified linear unit function, ensuring that it is set to zero if negative. The features extracted from the three inputs are represented by the terms $f\left({x}_{i}\right)$, $f\left({x}_{j}\right)$, and $f\left({x}_{k}\right)$.
4 Methodology
Driven by the challenges of unstable training and distribution shift in traditional RL alignment methods, we propose the TI-DPO framework. Our key innovation lies in integrating the token-importance weight and triplet loss in the TI-DPO framework, attaining a more optimal equilibrium balance between alignment precision and generative diversity.
4.1 Token-Level MDP for LLM Preference Alignment
To address the challenges of the sequential and auto-regressive nature of text generation, a token-level Markov Decision Process (MDP) is introduced, which incorporates the notion of token significance to improve the alignment of each token selection with human preferences. This concept is defined through a tuple denoted as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho_{0})$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}$ is a deterministic transition model among tokens. Here $r$ stands reward model associated with each token, and $\rho_{0}$ indicates the initial state distribution. The initial state is $s_{0}=[x]$, which is simply the input prompt. At each step $t$ of the generation process, the state $s_{t}=[x,y^{<t}]\in\mathcal{S}$ consists of input prompt $x=[x_{1},x_{2},\ldots,x_{m}]$, where $m$ is the count of token, and $t-1$ generated tokens $y^{<t}=[y^{1},y^{2},\ldots,y^{t-1}]$. At each time step $t$, the action $a_{t}=y^{t}$ corresponds to the selection of subsequent tokens for generation. The initial state distribution $\rho_{0}$ represents the distribution over input prompts $x$.
4.2 Gradient Attribution for Token Importance
Building on the token-level MDP framework, we now formalize the calculation of importance weights $w_{t}$. Inspired by the attribution-based rationale extraction from Ballout et al. [40], our approach quantifies token importance through gradient sensitivity analysis, ensuring that critical tokens in human-preferred responses drive the policy optimization process.
Assume that each token $x_{i}$ is transformed into an embedding vector $e_{i}$, forming an embedding sequence $E=[e_{1},e_{2},\cdots,e_{m}]$. We compute the gradient of reward model’s output concerning $e_{i}$ to measure how sensitive the reward is to changes in the token. For each reward $r_{\phi}(s_{j},a_{j})$, the gradient corresponding to $e_{i}$ is $\nabla_{e_{i}}r_{\phi}(s_{j},a_{j})$. Similar to previous work [40], the average gradient for embedding $e_{i}$ is given by:
$\bar{G}_{i}=\frac{1}{t}\sum_{j=1}^{t}\nabla_{e_{i}}r_{\phi}(s_{j},a_{j}).$ (6)
Since $\pi_{\text{ref}}$ is fixed, there is $\nabla_{e_{i}}\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}=\nabla_{e_{i}}\log\pi_{\theta}(y^{t}|x,y^{<t})$. For each token $x_{i}$ in the input or generated sequence, we rewrite (6) with policy log-probability ratio as:
$\bar{G}_{i}=\frac{1}{t}\sum_{j=1}^{t}\nabla_{e_{i}}\log\pi_{\theta}(y^{j}|x,y^{<j}).$ (7)
Although the average operation may weaken the long-range dependence, the generation of each token $y^{t}$ only only depends on the prefix $y^{<t}$, so the gradient $\nabla_{e_{i}}\log\pi_{\theta}(y^{j}|x,y^{<j})$ has implicitly included the historical dependence. We take the average of all time steps to balance the local and global importance. Denote the importance-score for the $i$-th token as $I_{i}$. Following the attribution method of Ballout et al. [40], we calculate $I_{i}$ using the $L_{1}$ norm of gradient:
$I_{i}=\left\|\bar{G}_{i}\right\|_{1}=\sum_{k}\left|\bar{G}_{i}[k]\right|.$ (8)
Here index $k$ corresponds to the components within the average gradient vector $\bar{G}_{i}$ [40]. This gradient-based importance guidance addresses the key limitations of previous work. Gradient attribution provides a data-driven measure of token relevance, which can adapt to the subtle semantics of human preferences and achieve fine-grained control of key tokens in the model generation process.
Then we normalize $I_{i}$ to derive the importance-guided token-wise weight $w_{t}\in[0,1]$. Let $\mathcal{T}_{t}$ be the set of tokens in $s_{t}$. To avoid division by zero, we assume that at least two tokens have distinct importance scores in $\mathcal{T}_{t}$. The normalized weight for token $i$ is:
$w_{t}=\frac{I_{i}-\min(\mathcal{I}_{t})}{\max(\mathcal{I}_{t})-\min(\mathcal{I}_{t})},$ (9)
where $\mathcal{I}_{t}=\{I_{j}|j\in\mathcal{T}_{t}\}$ is the set of importance scores for tokens. Normalize to $[0,1]$ to maintain consistent weight measurements across diverse samples. The BT-model that $y_{\text{w}}$ is favored compared with $y_{\text{l}}$ is:
$p^{*}(y_{\text{w}}\succ y_{\text{l}})=\frac{\exp\left(\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\cdot r_{\phi}(s_{t}^{\text{w}},a_{t}^{\text{w}})\right)}{\exp\left(\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\cdot r_{\phi}(s_{t}^{\text{w}},a_{t}^{\text{w}})\right)+\exp\left(\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\cdot r_{\phi}(s_{t}^{\text{l}},a_{t}^{\text{l}})\right)}.$ (10)
Here, $w_{t}^{\text{w}}$ and $w_{t}^{\text{l}}$ are the importance weights for tokens in the preferred and less-preferred tokens, $T_{\text{w}}$ and $T_{\text{l}}$ are the lengths of $y_{\text{w}}$ and $y_{\text{l}}$ respectively. Then with $r_{\phi}(s_{t},a_{t})=\beta\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}$ in DPO, we can derive the expression for BT model:
$p^{*}(y_{\text{w}}\succ y_{\text{l}})=\sigma(\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})),$ (11)
where $\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})$ can be denoted as:
$\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})=\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\log\frac{\pi_{\theta}\left(y^{t}_{\text{w}}\mid x,y_{\text{w}}^{<t}\right)}{\pi_{\text{ref }}\left(y^{t}_{\text{w}}\mid x,y_{\text{w}}^{<t}\right)}-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\log\frac{\pi_{\theta}\left(y^{t}_{\text{l}}\mid x,y_{\text{l}}^{<t}\right)}{\pi_{\text{ref }}(y^{t}_{\text{l}}\mid x,y_{\text{l}}^{<t})}.$ (12)
Therefore, we obtain the weighted token-level DPO base loss as [20]:
${\mathcal{L}_{\text{DPO-w }}=-\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\left[\log\sigma\left(\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})\right)\right]}.$ (13)
4.3 TI-DPO Objective and Theoretical Analysis
DPO defines the reward model as the log-likelihood ratio between $\pi_{\theta}$ and $\pi_{\text{ref}}$, eliminating the need for an explicit reward function. However, TI-DPO extends this by incorporating triplet loss, which explicitly guides outputs $y$ to align with preferred responses $y_{\text{w}}$ and diverge from non-preferred $y_{\text{l}}$. The dataset $D=\{(x,y_{\text{w}},y_{\text{l}})\}$ enables learning to pull responses toward high-quality standards and reject low-quality alternatives. Building on the triplet loss definition in Eq.(5), we obtain the triplet objective with policy model as follows:
$\mathcal{L}_{\text{triplet}}=$ $\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\Big{[}\max(0,\ \underbrace{\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}}_{\text{Bring }y\text{ closer to }y_{\text{w}}}$ (14)
$-\underbrace{\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}}_{\text{Push }y\text{ away }y_{\text{l}}}+\alpha)\Big{]}_{+}.$
[…]
$\mathcal{L}_{\text{TI-DPO}}=\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}},$ (15)
where $\gamma$ is a hyperparameter. This combined loss guides the policy $\pi_{\theta}$ to learn to assign similar scores to $y_{\text{w}}$, while emphasizing the importance of tokens in $y_{\text{w}}$ through the weighted reward function. In Appendix A.3, we have given the proof of gradient $\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$, which is used to update $\theta$ during training. Denote $\Delta r_{\text{global}}=\log\frac{\pi_{\theta}\left(y_{\text{w}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{w}}|x\right)}-\log\frac{\pi_{\theta}\left(y_{\text{l}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{l}}|x\right)}$ from Eq.(4). For simplicity, we abbreviate $\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})$ as $\Delta r_{\text{token}}$. To show the superiority of our TI-DPO loss compared to DPO loss, we first introduce the following lemma:
Lemma 4.1
There is $\alpha>1$ s.t. $\forall(x,y_{\text{w}},y_{\text{l}})\sim D$, it is satisfied that:
$\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right].$ (16)
The proof of Lemma 4.1 is detailed in Appendix A.1. According to Eq.(13), TI-DPO dynamically allocates token importance weights through gradient attribution, enabling the model to focus on the generation outputs with more importance weights on preference alignment. The following theorem strictly proves the theoretical advantages of this improvement at the loss function level. Under the condition of Lemma 4.1, the total loss of TI-DPO will be significantly lower than the original DPO loss:
Theorem 4.2 (Tighter Loss Bound)
If $\exists\alpha>1$ s.t. $\forall(x,y_{\text{w}},y_{\text{l}})\sim D$ satisfying $\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right]$, then TI-DPO’s total loss is strictly lower than DPO’s, i.e.,
$\mathcal{L}_{\text{TI-DPO}}\leq\mathcal{L}_{\text{DPO}}-\beta\Delta_{\text{triplet}},$ (17)
where $\Delta_{\text{triplet}}$ is a positive coefficient related to triplet loss and $\alpha$.
The proof is shown in Appendix A.2. Theorem 4.2 indicates that the total loss of TI-DPO is strictly less than the original DPO loss by an additional positive term determined jointly by the token-level weight advantage and the triplet coefficient. The strict superiority of $\mathcal{L}_{\text{TI-DPO}}$ provides a mathematical guarantee for the more stable training process of TI-DPO in practical tasks, verifying its essential improvement compared to the traditional DPO from a theoretical perspective. The implementation of TI-DPO is shown in Algorithm 1 as follows.
Algorithm 1 TI-DPO
1:Input: Dataset $\mathcal{D}=\{(x,y_{\text{w}},y_{\text{l}})\}$, hyperparameter $\beta$ and $\alpha$, the length of preferred responses $T_{\text{w}}$, the length of non-preferred responses $T_{\text{l}}$, reference model $\pi_{\text{ref}}$, policy model $\pi_{\theta}$.
2:Initialize: $\pi_{\theta}\leftarrow\pi_{\text{ref}}$
3:for each epoch do
4:     Sample batch $\{(x,y_{\text{w}},y_{\text{l}})\}\sim\mathcal{D}$.
5:     Compute policy model $\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})$ and $\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})$.
6:     Compute reference model $\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})$ and $\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})$.
7:     Calculate importance weights for preferred responses:
8:     $w_{t}^{\text{w}}=\text{Normalize}(\|\frac{1}{T_{\text{w}}}\sum_{t=1}^{T_{\text{w}}}\nabla_{e_{t}}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\|_{1}).$
9:     Calculate importance weights for non-preferred responses:
10:     $w_{t}^{\text{l}}=\text{Normalize}(\|\frac{1}{T_{\text{l}}}\sum_{t=1}^{T_{\text{l}}}\nabla_{e_{t}}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}\|_{1}).$
11:     Compute weighted DPO log-ratio:
12:     $\Delta r_{\text{token}}=\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}-\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$
13:     Calculate the output $y^{t}$: $y^{t}\sim\pi_{\theta}(y^{t-1}|x,y^{<t-1})$.
14:     Compute triplet log-ratio:
15:     $\Delta r_{\text{triplet}}=\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}-\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}$.
16:     Compute weighted DPO loss: $\mathcal{L}_{\text{DPO-w}}=-\log\sigma(\beta\Delta r_{\text{token}})$.
17:     Compute triplet loss: $\mathcal{L}_{\text{triplet}}=\max\left(0,\Delta r_{\text{triplet}}+\alpha\right)$.
18:     Aggregate losses: $\mathcal{L}_{\text{TI-DPO}}\leftarrow\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}}$.
19:     Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$.
20:end for
21:Output: $\pi_{\theta}$
5 Experiments
This section conducts extensive experiments to evaluate our proposed TI-DPO. Our experiments show the effectiveness of token-importance guided weight and analyze the performance on model alignment compared with other works.
5.1 Experimental Settings
Dataset and base settings: The following datasets are employed across various tasks to evaluate our approach. MMLU [43] is used to assess the model’s knowledge breadth, covering multiple-choice tasks in 57 disciplines. GSM8K [44] and MATH [45] focus on mathematical reasoning, including primary school math application problems and advanced math problems, respectively. MT-Bench [47] and its enhanced version Arena-Hard are judged by GPT-4, examining the conversation ability and complex instruction processing. Additionally, TruthfulQA [49] detects the authenticity of the model’s answers through adversarial questions. GPQA [50] evaluates graduate-level scientific question-answering with 448 multi-disciplinary multiple-choice questions, and IFEval [51] measures instruction-following accuracy through  500 verifiable tasks. HumanEval [52] examines the code generation ability with the pass@1 indicator.
Comparative algorithm: We compared the TI-DPO with baseline alignment methods such as SFT, DPO [14], IPO[53], KTO [33], SimPO [19], TDPO [17], CPO [16], TPO [25], and GRPO [28]. In different alignment tasks, we selected a base model that balances performance and efficiency, comprehensively considering the model parameter scale and adaptability to guarantee the fairness and efficacy of the comparative experiment. Specifically, we select three models (Llama-3.2-3B [54], Llama-3.1-8B [54], Mistral-7B-v0.3 [55]) as baselines.
5.2 Performance Comparison
As shown in Figure 1, we conduct an analysis of the performance for TI-DPO and baseline methods across training steps on the TruthfulQA (reliability assessment) and IFEval (instruction-following) tasks with Llama-3.1-8B model. In the TruthfulQA benchmark (Figure 1a), TI-DPO demonstrates a steady improvement in accuracy as training steps increase, surpassing all baselines by the final epoch. For the IFEval task (Figure 1b), TI-DPO also shows a dominant performance trend. This highlights TI-DPO’s effectiveness in learning through token-level importance weighting and triplet loss, which explicitly guides the model to avoid generating misleading content.
(a) TruthfulQA Performance vs. Training Steps
(b) IFEval Performance vs. Training Steps
Figure 1: Accuracy trends with training steps for different methods on TruthfulQA and IFEval tasks on LLaMA-3.1-8B. The performance comparisons of SFT, DPO, IPO, KTO, SimPO, TDPO, CPO, TPO, GRPO, and TI-DPO are illustrated.
As shown in Figure 2, TI-DPO exhibits significantly better performance in Reasoning, Instruction-Following, and Reliability dimensions compared to the corresponding instruction variants with each base instruct. Here, the base instruct refers to the foundational instruction-tuned models (Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3), serving as the baseline for comparing the effectiveness of TI-DPO and other fine-tuning methods. The scores of TI-DPO in other aspects are roughly equal or slightly higher than others. Table 1 presents average scores of each fine-tuning method across three base models, clearly demonstrating our method’s advantages in general tasks and specific scenarios. The specific score comparison table under the three base model is placed in Appendix B.
Figure 2: Multi-dimensional normalized score of TI-DPO compared with other base instruction models across categories.
Table 1: Average scores of each fine-tuning method across three base models
Method MMLU GSM8K GPQA HumanEval TruthfulQA IFEval Avg
SFT 64.0 68.0 22.7 59.3 55.5 70.5 56.7
DPO 65.3 69.3 24.0 61.0 56.7 70.0 57.7
IPO 63.0 65.3 20.3 57.3 52.7 66.7 54.2
KTO 66.3 70.3 25.3 62.0 57.7 70.5 58.7
SimPO 63.5 64.7 21.8 58.2 54.2 64.7 54.5
TDPO 65.0 68.2 23.5 60.3 56.3 68.5 57.0
CPO 67.3 70.7 26.0 62.8 58.3 71.3 59.4
TPO 68.3 72.7 27.7 63.7 59.0 72.7 60.7
GRPO 70.7 75.7 28.0 64.3 59.9 74.0 62.1
TI-DPO 70.0 73.0 26.0 67.0 62.0 75.7 62.3
5.3 Ablation Experiment
Table 2: Ablation study scores: the full TI-DPO vs. base instruction model (Llama-3.2-3B-Instruct) with other weight and triplet Conditions
Method General Math Reasoning Code Instr-Follow Reliability
Base Instruct (Baseline) 0.634 0.777 0.266 0.280 0.515 0.768
Full Method (TI-DPO) 0.654 0.807 0.346 0.330 0.635 0.868
No Triplet Loss 0.640 0.790 0.320 0.310 0.605 0.830
Uniform Weight 0.640 0.782 0.305 0.290 0.580 0.800
Random Weight 0.637 0.778 0.280 0.285 0.550 0.780
In the ablation experiment section, we conduct an ablation analysis on the impact of the "importance guidance" mechanism and the "triplet loss" on the model performance. It focuses on tasks where the impact of the expected importance mechanism such as GSM8K [44] for mathematical reasoning and TruthfulQA [49] for reliability is significant.
In order to verify the improvement effect of various weight and triplet conditions, Table 2 shows the ablation experiment with Llama-3.2-3B-Instruct as baseline. The experimental design compares TI-DPO (full method), baseline with randomized weights, baseline with uniform importance weights but ignoring the importance differences of tokens, as well as different configurations with or without the triplet loss. It can be observed that the ordered design of weight (from Baseline $\rightarrow$ Random weight $\rightarrow$ Uniform Weight $\rightarrow$ Full Method) leads to a gradually enhanced performance in all aspects. The design of the triplet loss comprehensively improves the performance score, especially for Instruction-Following and Reliability.
5.4 Case Demo
We present a medical scenario case demo in Fig. 3 to illustrate how TI-DPO dynamically assigns token importance weights and leverages triplet loss to optimize model outputs. Given the user prompt "I have a headache, what should I do?", three responses are analyzed: preferred response (A), intermediate response (B), and non-preferred response (C). More details of this demo would be shown in Appendix B.2.
Figure 3: Case demo of responses to prompt "I have a headache, what should I do?". Left: Preferred case. Middle: Inermedia case. Right: Non-preferred case. The darker color indicates higher weight.
6 Conclusion
We introduce TI-DPO, an optimization framework that effectively bridges the alignment gap between LM and human value systems. By incorporating token-importance guidance and a triplet loss, TI-DPO effectively overcomes the limitations of traditional DPO methods, which often neglect individual token significance and are sensitive to judgment noise. Rigorous experiments across diverse benchmarks such as HumanEval, TruthfulQA, and IFEval show that TI-DPO consistently outperforms baselines like SFT, DPO, and their variants. In performance comparisons, TI-DPO achieves the highest accuracy on TruthfulQA and IFEval tasks, demonstrating its superiority in fine-grained preference alignment. And we compare TI-DPO with three base instruction models across multiple task dimensions with radar graphs. Additionally, ablation experiments confirm the necessity of its token-importance guidance and triplet loss components. Visualized weight distributions in the appendix reveal that TI-DPO focuses on semantically critical tokens, such as safety terms of medical responses that we show in the case study, while theoretical analysis proves it achieves a tighter loss bound than DPO, ensuring optimization stability. These experiments highlight the good performance of TI-DPO in improving the quality and reliability of LLM output, providing a powerful solution for preference optimization.
Appendix A Theoretical Proof
A.1 Proof of Lemma 4.1
Using the linear property of expectation, the expectation of the weighted reward difference between the preferred response $y_{\text{w}}$ and the non-preferred response $y_{\text{l}}$ can be decomposed as:
$\mathbb{E}[\Delta r_{\text{token}}]$ $=\mathbb{E}\left[\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})-\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})\right]$ (18)
$=\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]-\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})].$
where $T_{w}$ and $T_{l}$ are the token lengths of the preferred and non-preferred responses respectively, and $w_{t}^{w}$ and $w_{t}^{l}$ are the token-importance weights calculated dynamically through gradient attribution. For the token $t$ in the preferred response, since the weight $w_{t}^{w}$ is positively correlated with the reward $r_{\phi}(x,y_{w}^{t})$, i.e., $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\geq 0$, then we have
$\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]=\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})]+\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))$ (19)
Similarly, for $y_{\text{l}}$, the weight $w^{\text{l}}_{t}$ is negatively correlated with its reward $r_{\phi}(x,y_{\text{l}}^{t})$, i.e., $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\leq 0$. Then, we have:
$\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})]=\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})]+\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))$ (20)
Substitute the positive correlation condition $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\geq 0$ into (19), we obtain
$\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]\geq\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})],$ (21)
Similarly, with $\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))\leq 0$ for (20), we have:
$\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})]\leq\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})].$ (22)
Denote $\lambda=\min\big{\{}\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t},\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}\big{\}}>1$, we have
$\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})]\geq\lambda\mathbb{E}\left[\sum_{t=1}^{T_{\text{w}}}r_{\phi}(x,y_{\text{w}}^{t})\right]=\lambda\mathbb{E}[r_{\phi}(x,y_{\text{w}})],$ (23)
and
$\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})]\geq\lambda\mathbb{E}\left[\sum_{t=1}^{T_{\text{l}}}r_{\phi}(x,y_{\text{l}}^{t})\right]=\lambda\mathbb{E}[r_{\phi}(x,y_{\text{l}})].$ (24)
Substituting (21) and (22) into (18), there is
$\mathbb{E}[\Delta r_{\text{token}}]$ $\geq\lambda\mathbb{E}[r_{\phi}(x,y_{\text{w}})]+\sum_{t=1}^{T_{\text{w}}}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))-\left(\lambda\mathbb{E}[r_{\phi}(x,y_{\text{l}})]-\sum_{t=1}^{T_{\text{l}}}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|\right)$ (25)
$=\lambda\mathbb{E}[\Delta r_{\text{global}}]+\underbrace{\sum_{t=1}^{T_{\text{w}}}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))+\sum_{t=1}^{T_{\text{l}}}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|}_{C>0}.$
Let $C=\sum_{t=1}^{T}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))+\sum_{t=1}^{T}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|$, $C>0$ and define $\alpha=\lambda+\frac{C}{\mathbb{E}[\Delta r_{\text{global}}]}.$ Since the global reward of the preferred response $y_{\text{w}}$ is necessarily higher than that of the non-preferred response $y_{\text{l}}$, i.e., $\mathbb{E}[\Delta r_{\text{global}}]>0$, we have $\alpha>1$, and
$\mathbb{E}[\Delta r_{\text{token}}]\geq\alpha\cdot\mathbb{E}[\Delta r_{\text{global}}].$ (26)
Hence, we show the Lemma 4.1.
A.2 Proof of Theorem 4.2
Since the sigmoid function $\sigma(x)=1/(1+\exp(-x))$ is convex, using Jensen’s inequality to $\mathcal{L}_{\text{DPO-w}}$, we have
$\mathcal{L}_{\text{DPO-w}}=-\mathbb{E}\left[\log\sigma\left(\beta\Delta r_{\text{token}}\right)\right]\leq-\log\sigma\left(\beta\mathbb{E}\left[\Delta r_{\text{token}}\right]\right)$ (27)
According to the assumption $\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right]$ shown in Appendix 1, with the monotonicity of $\sigma(z)$, we have
$-\log\sigma\left(\beta\mathbb{E}\left[\Delta r_{\text{token}}\right]\right)\leq-\log\sigma\left(\beta\alpha\mathbb{E}\left[\Delta r_{\text{global}}\right]\right).$ (28)
Denote $z=\beta\mathbb{E}\left[\Delta r_{\text{global}}\right]>0$ and construct the function $g(\alpha)=\log(1+e^{-\alpha z})+(\alpha-1)z$. Since the monotonicity and $g(1)=\log(1+e^{-z})$, when $\alpha>1$, we have $g(\alpha)<g(1)$, i.e.,
$-\log\sigma(\alpha z)=\log(1+e^{-\alpha z})\leq\log(1+e^{-z})-(\alpha-1)z=-\log\sigma(z)-(\alpha-1)z$ (29)
Substituting $z=\beta\mathbb{E}\left[\Delta r_{\text{global}}\right]$ into (29), we have
$\mathcal{L}_{\text{DPO-w}}\leq\mathcal{L}_{\text{DPO}}-\beta(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ (30)
According to Eq.(14), when the optimization is close to being completed,
$\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}$ (31)
$-\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}\geq-\alpha$
and there exists a constant $\Delta_{\text{triplet}}$ such that $\mathcal{L}_{\text{triplet}}\leq\Delta_{\text{triplet}}$. Assuming $\gamma<\beta$ and combining the importance-weighted DPO loss and triplet loss, we have
$\mathcal{L}_{\text{TI-DPO}}$ $=\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}}$ (32)
$\leq\mathcal{L}_{\text{DPO}}-\beta(\alpha-1)\mathbb{E}[\Delta r_{\text{global}}]+\beta\Delta_{\text{triplet}}$
Let $\Delta_{\text{triplet}}=\frac{1}{2}(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ and substitute into (32), then we have
$\mathcal{L}_{\text{TI-DPO}}$ $\leq\mathcal{L}_{\text{DPO}}-\frac{1}{2}\beta(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ (33)
$\leq\mathcal{L}_{\text{DPO}}-\beta\Delta_{\text{triplet}}$
Hence, Theorem 4.2 holds.
A.3 Gradient Analysis of Loss
According to Eq.(13), we have $\mathcal{L}_{\text{DPO-w}}=-\log\sigma(\beta\Delta r_{\text{token}}),$ where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function with derivative $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$. Taking the gradient of Eq.(13), we have
$\nabla_{\theta}\mathcal{L}_{\text{DPO-w}}$ $=-\frac{1}{\sigma(\beta\Delta r_{\text{token}})}\cdot\sigma^{\prime}(\beta\Delta r_{\text{token}})\cdot\beta\nabla_{\theta}\Delta r_{\text{token}}$ (34)
$=-\beta(1-\sigma(\beta\Delta r_{\text{token}}))\nabla_{\theta}\Delta r_{\text{token}}.$
Here, we expand $\nabla_{\theta}\Delta r_{\text{token}}$:
$\nabla_{\theta}\Delta r_{\text{token}}=\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$ (35)
Since $\pi_{\text{ref}}$ is fixed, there is $\nabla_{\theta}\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}=\nabla_{\theta}\log\pi_{\theta}(y^{t}|x,y^{<t})=\frac{1}{\pi_{\theta}(y^{t}|x,y^{<t})}\nabla_{\theta}\pi_{\theta}(y^{t}|x,y^{<t}).$
Thus, $\nabla_{\theta}\mathcal{L}_{\text{DPO-w}}$ becomes:
$\nabla_{\theta}\mathcal{L}_{\text{DPO-w}}=-\beta(1-\sigma(\beta\Delta r_{\text{token}}))\left[\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})\right].$ (36)
As for the gradient of $\mathcal{L}_{\text{triplet}}=\mathbb{E}\left[\max\left(0,\Delta r_{\text{triplet}}+\alpha\right)\right]_{+}$, assuming $\Delta r_{\text{triplet}}+\alpha>0$, we can expand the gradient:
$\nabla_{\theta}\mathcal{L}_{\text{triplet}}=\nabla_{\theta}\sum_{t=1}^{T_{\text{w}}}\left\|d_{t}-b_{t}\right\|_{2}^{2}-\nabla_{\theta}\sum_{t=1}^{T_{\text{l}}}\left\|c_{t}-d_{t}\right\|_{2}^{2},$ (37)
where $b_{t}=\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}$, $c_{t}=\log\frac{\pi_{\theta}(y_{l}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}$, $d_{t}=\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}$ for simplicity. Differentiating the squared terms, we have
$\nabla_{\theta}\left\|d_{t}-b_{t}\right\|_{2}^{2}=2(d_{t}-b_{t})\left(\nabla_{\theta}d_{t}-\nabla_{\theta}b_{t}\right),$ (38)
and
$\nabla_{\theta}\left\|c_{t}-d_{t}\right\|_{2}^{2}=2(c_{t}-d_{t})\left(\nabla_{\theta}c_{t}-\nabla_{\theta}d_{t}\right).$ (39)
Then, we substitute the definitions of $b_{t},c_{t},d_{t}$ and use $\nabla_{\theta}\log\pi_{\text{ref}}=0$:
$\nabla_{\theta}b_{t}=\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t}),$ (40a)
$\nabla_{\theta}c_{t}=\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t}),$ (40b)
$\nabla_{\theta}d_{t}=\nabla_{\theta}\log\pi_{\theta}(y^{t}|x,y^{<t}).$ (40c)
Thus, $\nabla_{\theta}\mathcal{L}_{\text{triplet}}$ is:
$\nabla_{\theta}\mathcal{L}_{\text{triplet}}=2\sum_{t=1}^{T_{\text{w}}}(d_{t}-b_{t})\left(\nabla_{\theta}d_{t}-\nabla_{\theta}b_{t}\right)-2\sum_{t=1}^{T_{\text{l}}}(c_{t}-d_{t})\left(\nabla_{\theta}c_{t}-\nabla_{\theta}d_{t}\right).$ (41)
Substituting the derived gradients:
$\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}=$ $-\beta(1-\sigma(\beta\Delta r_{\text{token}}))\left[\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})\right]$ (42)
$+2\gamma\left[\sum_{t=1}^{T_{\text{w}}}(d_{t}-b_{t})(\nabla_{\theta}d_{t}-\nabla_{\theta}b_{t})-\sum_{t=1}^{T_{\text{l}}}(c_{t}-d_{t})(\nabla_{\theta}c_{t}-\nabla_{\theta}d_{t})\right]$
Appendix B Additional Experimental Results
We present some additional explanations and experimental results.
B.1 Distribution of Weights
Figure 4 visualizes the distribution of importance weights assigned to tokens in different tasks by TI-DPO with histograms and box plots, intuitively explaining how the importance mechanism dynamically focuses on key tokens according to task characteristics. The y-axis of the histogram is token frequency, which represents the number of occurrences of tokens with varying importance. The box plots in the upper right corner are used to visualize the distribution of token frequencies for different weights, where the red line represents the median, reflecting the concentration of token frequencies. In the GSM8K and GPQA datasets, with only a few key symbols crucial for answering, the importance weights mostly concentrate in the interval of $[0.2,0.5]$ approximately. For the TruthfulQA and IFEval datasets, token distribution is concentrated in the $[0.6,0.8]$ weight interval. The MMLU and HumanEval datasets cover a wide range of content, and the model assigns diverse weights to various tokens.
Figure 4: Distribution patterns of gradient-based token importance weights in six benchmark tasks (GSM8K, TruthfulQA, MMLU, GPQA, HumanEval, IFEval)
B.2 Case Demo
In the TI-DPO framework, response A corresponds to the preferred response $y_{\text{w}}$ in the dataset, which represents the high-quality, human-preferred output aligned with safety, professionalism, and correctness (e.g., "seek medical attention promptly" in the medical case). These tokens are assigned high importance weights to prioritize critical elements in human judgments. As show in Table 5, critical tokens like "seek" (0.93), "medical" (0.87), "attention" (0.85), "promptly" (1.00), "avoid" (0.92), and "self-medicating" (0.89) receive high importance weights, reflecting their role in ensuring safety and compliance with medical standards. These weights act as a "spotlight" to prioritize tokens that most influence human judgments, such as emergency actions and avoidance of self-treatment.
Response C in Table 5 corresponds to the less preferred response $y_{\text{l}}$, representing low-quality or risky outputs that deviate from human preferences (e.g., "take painkillers casually" in the example). The model includes high-risk tokens like "painkillers" (0.91), "casually" (1.00), and "fine" (0.97), which receive peak weights due to their potential to mislead users into unsafe self-medication. TI-DPO’s gradient-based attribution mechanism identifies these tokens as critical for preference misalignment, suppressing their influence during generation.
In Table 5, response B represents an intermediate generated response (e.g., "get more rest… consult a doctor"), which is neither the top preferred nor the worst case. In the triplet loss structure, B acts as an anchor that is guided to approach $y_{\text{w}}$ (A) and distance from $y_{\text{l}}$ (C). Key tokens like "worsen" (0.88), "should" (0.82), "consult" (0.90), and "doctor" (0.95) have elevated weights but are less intense than those in A, indicating their secondary importance in guiding less urgent but still reasonable advice. By incorporating B, TI-DPO promotes more nuanced optimization, where intermediate outputs are refined to better match human preferences through token-level importance weights and triplet constraints.
Table 3: Token Importance
Assignment of A
Token Weight
Based 0.05
on 0.05
your 0.05
symptoms 0.18
it 0.03
is 0.03
recommended 0.20
that 0.02
you 0.02
seek 0.93
medical 0.87
attention 0.85
promptly 1.00
and 0.03
avoid 0.92
self-medicating 0.89
Table 4: Token Importance
Assignment of B
Token Weight
According 0.04
to 0.04
your 0.04
description 0.07
it 0.03
is 0.03
advised 0.13
to 0.02
get 0.06
more 0.06
rest 0.11
symptoms 0.18
worsen 0.88
you 0.03
should 0.82
consult 0.90
doctor 0.95
Table 5: Token Importance
Assignment of C
Token Weight
Don’t 0.21
worry 0.18
you 0.04
can 0.04
just 0.09
take 0.09
some 0.05
painkillers 0.91
casually 1.00
it 0.02
should 0.06
be 0.03
fine 0.97
B.3 Other Tables
This subsection presents evaluation results of TI-DPO on three different base models (LLaMA-3.2-3B, LLaMA-3.1-8B, Mistral-7B-v0.3), comparing it with baseline methods like SFT, DPO, and other variants to validate its effectiveness and robustness across model scales and tasks.
Table 6 presents the evaluation results of TI-DPO on the LLaMA-3.2-3B model, a lightweight 3B-parameter model, showing that TI-DPO achieves notable scores of 68.0 in HumanEval and 82.0 in IFEval, outperforming baselines like DPO (62.0, 78.0) and SFT (61.0, 77.4) significantly. Table 7 evaluates TI-DPO on the LLaMA-3.1-8B model (8B parameters), where it excels with an IFEval score of 86.0, surpassing GRPO (85.0), and achieves 80.0 in HumanEval and 63.0 in TruthfulQA, outperforming DPO (74.0, 58.0) and GRPO (78.0, 62.0); with an average score of 71.1, it closely matches the best baseline, validating its capability to handle complex instructions and improve generative reliability on medium-scale models. In Table 8, TI-DPO achieves 66.0 in TruthfulQA and 59.0 in IFEval, significantly higher than DPO (60.0, 50.0), and surpasses GRPO in HumanEval (53.0 vs. 51.0).
Table 6: LLaMA-3.2-3B evaluation
Method MMLU GSM8K GPQA HumanEval TruthfulQA IFEval Avg
SFT 63.0 78.0 33.0 61.0 51.0 77.4 60.6
DPO 64.0 79.0 34.0 62.0 52.0 78.0 61.5
IPO 62.0 76.0 31.0 59.0 49.0 76.0 58.8
KTO 65.0 80.0 35.0 63.0 53.0 78.5 62.4
SimPO 64.0 78.0 33.5 61.5 51.5 74.0 60.4
TDPO 64.5 78.5 34.0 62.0 52.0 76.5 61.2
CPO 66.0 79.5 35.5 63.5 53.5 79.0 62.8
TPO 67.0 82.0 39.0 64.0 54.0 80.0 64.3
GRPO 69.0 85.0 38.0 63.8 53.8 81.0 65.1
TI-DPO 68.0 81.0 34.5 68.0 57.0 82.0 65.1
Table 7: LLaMA-3.1-8B evaluation
Method MMLU GSM8K GPQA HumanEval TruthfulQA IFEval Avg
SFT 69.0 84.0 30.0 72.0 56.0 80.0 65.2
DPO 70.0 85.0 32.0 74.0 58.0 82.0 66.8
IPO 68.0 80.0 27.0 70.0 53.0 77.0 62.5
KTO 71.0 86.0 34.0 75.0 59.0 83.0 68.0
SimPO 68.5 78.0 28.0 71.0 54.0 75.0 62.4
TDPO 69.5 83.0 31.0 73.0 57.0 81.0 65.8
CPO 72.0 86.5 35.0 76.0 60.0 84.0 68.9
TPO 73.0 88.0 36.0 77.0 61.0 85.0 70.0
GRPO 75.0 90.0 37.0 78.0 62.0 85.0 71.2
TI-DPO 74.0 89.0 34.5 80.0 63.0 86.0 71.1
Table 8: Mistral-7B-v0.3 evaluation
Method MMLU GSM8K GPQA HumanEval TruthfulQA IFEval Avg
SFT 60.0 42.0 5.0 45.0 59.5 54.0 44.2
DPO 62.0 44.0 6.0 47.0 60.0 50.0 44.8
IPO 59.0 40.0 3.0 43.0 56.0 47.0 41.3
KTO 63.0 45.0 7.0 48.0 61.0 50.0 45.7
SimPO 58.0 38.0 4.0 42.0 57.0 45.0 40.7
TDPO 61.0 43.0 5.5 46.0 60.0 48.0 43.9
CPO 64.0 46.0 7.5 49.0 61.5 51.0 46.5
TPO 65.0 48.0 8.0 50.0 62.0 53.0 47.7
GRPO 68.0 52.0 9.0 51.0 64.0 56.0 50.0
TI-DPO 66.0 47.0 7.0 53.0 66.0 59.0 49.7